### v1

@@ 1 Introduction
Our results reveal that VLMs leak identities even when modalities are aligned using anonymized datasets. This is persistence regardless of the subject’s context, showing interesting generalization capabilities. Moreover, widely adopted anonymization techniques, i.e. blurring, do not prevent privacy leakages. These results are surprising as they suggest that fine-tuning aligns unimodal (and sensible) knowledge even if not explicitly present in the training set. As a consequence, they demonstrate the need to go beyond simple data processing as a privacy protection strategies for VLMs, developing stronger strategies with more guarantees to the users. As VLMs are widely adopted, the key message of our findings is to increase our ethical awareness as a community for responsible development (and sharing) of such powerful models.
@@ 2.2 Function
As the name suggests, the objective is to embed both image and text in the same latent space, in a way that they can be compared by computing pairwise similarity. This is a core part for retrieval tasks. The matching is performed though a simple mechanism, which involves only $\boldsymbol{z}_{v}$ and $\boldsymbol{z}_{t}$. Once mapped in the same latent space, the similarity between the text and image can be estimated $s(\boldsymbol{z}_{v},\boldsymbol{z}_{v})$.
@@ 3.1 Experimental setting
Figure 3: Example of the picture manipulation we evaluated. (a) Original picture. (b) Background replaced with a random landscape background. (c) Background replace with white to force the model focus towards the subject. (d) Face blur to question its effectiveness in preventing leakages.
The first step is to collect a dataset of people whose images are contained in pre-training datasets, e.g., LAION [37]. The choice goes to publicly available pictures of celebrities, following a similar approach of [34] and [14]. We collect pictures of celebrities from publicly available sources, as no datasets with identity-face association are currently openly available. We selected 500 (Appendix  0.D) celebrities and collected 50 pictures each, summing up to 25k images for our experiments.
To remove potential sources of errors, we excluded pictures containing written names, which could guide VLMs to correctly recognizing the person only from the OCR capability. To do so we relied on EasyOCR [1] to detect text in the pictures, excluding only those containing a sub-string of the name with a minimum length of 5 characters. In total, only 1171 images have been removed.
@@
Generative VLMs are sensitive to the specific prompt used to condition their outputs. To avoid that our findings are limited by specific prompt choices, we studied models’ behavior when conditioned on different prompt categories, and their rephrasing. We identified five prompts representing increasing level of details required from the model:
@@
Interestingly, while recent research has analyzed identity recognition using contrastive VLMs [14], the privacy risks and the memorization of individuals’ identities remain unexplored in the context of generative VLMs. We opt for 5 VLMs pre-trained on potentially private data [37] and fine-tuned with few to none private data:
@@ 3.2 Do VLMs leak names?
Model $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$ BLIP-2flan-t5-xl [20] $0.95\pm 0.12$ $1.01\pm 0.88$ $6.94\pm 0.02$ $4.49\pm 6.2$ $7.21\pm 0.0$ BLIP-2opt [20] $1.43\pm 0.49$ $2.85\pm 3.41$ $5.24\pm 3.75$ $3.47\pm 3.0$ $4.6\pm 4.99$ LLaVA-1.57B [24] $0.12\pm 0.0$ $0.06\pm 0.0$ $4.14\pm 8.99$ $3.59\pm 9.69$ $9.56\pm 0.31$ LLaVA-1.6mixtral-7B [24] $0.35\pm 0.05$ $0.18\pm 0.0$ $7.12\pm 1.74$ $1.5\pm 2.23$ $6.83\pm 1.02$ PaliGemma3b-mix-224 [5] $1.12\pm 1.98$ $2.79\pm 2.57$ $7.81\pm 0.73$ $6.18\pm 2.08$ $7.66\pm 0.7$
@@
Looking at the results we first see that name leakages (or identity recognition) rate is not negligible, with a peak rate of 9% ($\sim$2250) of pictures whose subject has been identified by LLaVA-1.57B [24]. Second, all models leak more when directly asked to generate a person name, but can still leak correct identities even when prompted to generically describe the picture. For instance, the average leakage rate under $P_{2}$ (“Who is the person in the picture”) is $\sim$6%, with a peak of 7.81% with PaliGemma3b-mix-224 [5], while with $P_{3}$ (“Describe the celebrity in the picture.”), the average leakage rate is 3.4%. Looking at the most generic prompt $P_{0}$ (“Describe the picture.”), we see a decrease in leakages, but still experience a not indifferent amount of them with BLIP-2flan-t5-xl [20] (1.43% or 357 leaked images) and with PaliGemma3b-mix-224 [5] (1.12% or 280 leakaged images). This not only implies that identity recognition can be a threat, but also that it could happen in an unpredictable way, i.e., even when the model is not explicitly asked to perform it. We underline that this happens even when models are fine-tuned on anonymized data. In fact we don’t see any significant reduction in leakage for VLMs finetuned with anonymoized data compared to those finetuned with partially anonymzied data (BLIP-2). This clearly indicates that data anonymization does not address leakage if parts of the model (e.g., visual/textual backbones) are already exposed to private information. Additionally, we note that some models (i.e., LLaVA-1.57B [24] and BLIP-2opt [20]) are more susceptible to prompt rephrasing, as the high standard deviations show.
In all the remaining experiments we use the 5 prompts reported in Sec. 3.1.2 which overall 8 obtained results closer to the mean.
@@ 3.3 How does background influence leakages?
For instance, it is easier to identify Cristiano Ronaldo in a football field, rather than Cristiano Ronaldo alone, given the positive correlation between the two elements. To disentangle this correlation and focus primarily on the subject, we isolate the celebrities in each picture and change the background. To do so we used Grounding Dino [26] to first detect the class “person” in the picture, identifying their bounding boxes. After this, we forwarded each box to SAM [19] to obtain each subject’s segmentation masks. Finally we extracted the person’s masks and blend it with different backgrounds as shown in Figure 3 (b) and (c). We consider two cases of background choice.
@@
Results in Table 4 confirms that the reducing context distractions (as in landscape backgrounds) leads to higher leakages compared to cases where the background is a scene. For instance, LLaVA-1.6mixtral-7B [24] goes from 5.18% on $P_{4}$ with landscape background to 6.41% with white background. This highlights that landscape background can introduce novel patterns that weaken model’s focus on the subject, degrading their performance in recognizing identities. However, the overall performance remain in the same range, demonstrating that in both settings the model is able to focus on the subject and recognize their identity. Again, we underline how robust the identity recognition is, as not only all the models are able to recognize people from images that could have been precedently memorized, but instead they are able to recall the identity association even in different contexts.
@@ 3.4 Can blurring prevent privacy leakages?
This outcome suggests that blurring, without compromising the content of an image, may not an adequate technique for anonymization. A malicious attacker could use, for instance, Google Maps to recognize people whose pictures have been inadvertently included in a pretraining dataset.
3.4.2 Blurring on replaced background.
@@ 3.5 Statistical analysis on leaked celebrities and discussion
To further complement the experiments, we conduct a statistical analysis on the leaked celebrities, trying to extract insights that may suggest the origin of this behavior.
@@
We computed the leakage rate per celebrity (Figure 4, showing that more famous celebrities seem to be identified with more ease. Then we plot the distribution of the leakages per celebrity 5(a), which highlights that while a discrete portion of celebrities have been rarely identified, another portion of them have been easily identified throughout all the settings.
@@
As major continual learning works [23, 2, 43] emphasize the impact of catastrophic forgetting after additional sessions of training, face-identity preservation after fine-tuning is unexpected behaviour. This may be partially explained by the phenomenon known as memorization [11, 28]. When abnormal or complex patterns, such as the names of people, are presented to a model, it tends to memorize these specific instances rather than broadly learning general concepts. This phenomenon suggests that face-identity associations might be stored within a few neurons of the model as distinct activation patterns. These stored patterns can potentially be retrieved with some degree of context generalization, meaning that even slight contextual clues could trigger the recall of specific identities. This behavior raises significant privacy concerns, as it implies that sensitive information could be inadvertently exposed. Therefore, this discovery opens a promising direction for future research aimed at preventing such privacy leakages from occurring in generative Visual-Language Models (VLMs). Researchers will need to develop methods to ensure that models learn generalizable concepts without memorizing sensitive or personally identifiable information, thereby safeguarding user privacy.
4 Related Works
Vision Language Models
combine visual and textual data to solve tasks where both modalities are involved, such as Visual Question Answering [3], Image Captioning [15], Image-Text Retrieval [6] and Visual Reasoning [40]. Depending on the loss function and specific task they aim to address, VLMs are typically categorized into two main families: contrastive VLMs and generative VLMs. Contrastive VLMs are trained to assess the similarity between an image and a text. This paradigm includes widely adopted architectures like CLIP [34] and ALIGN [16] (employing batch-wise contrastive learning) and SigLIP [46] (using sigmoid based loss function).
On the other hand, generative VLMs are more powerful, as trained to generate textual descriptions or responses based on visual inputs. This family includes a diverse range of models such as BLIP [20], LLaVA [25], MiniGPT-4 [48], PaliGemma [5], and X-VLM [45]. Each of these models approaches the task with different nuances. For instance BLIP [20] is the first model which propose to leverage frozen modalities and focus on the training of the fusion model. Instead, LLaVA [25] stresses the relevance of more structured data, i.e., instruction data, and removes the fusion module of BLIP [20], preferring a fine-tuning of the entire architecture. In this work, we do not propose a training paradigm or architecture but show that these architectures, may leak private information, a problem future research should account for.
Privacy-Preserving AI
seeks to protect individual data while leveraging its value for AI applications. In this direction, differential privacy [12], allows data holders to share statistical analysis while limiting what can be inferred about specific individuals by injecting calibrated noise into statistical computations. For instance, in a healthcare database, differential privacy can be applied by adding random noise to each person’s age value so that the average age of the patients can be calculated without revealing any specific individual’s age. Papernot et al [33] employs differential privacy to train models without exposing individual data points. Differently, federated learning [29] trains models on decentralized devices, keeping data local to enhance privacy. Alternatively, data anonymization protects user privacy while retaining dataset utility [36, 27, 21]. In recognition application a simple data anonymization technique may involve blurring the face of individuals and other private information, to avoid re-identification [22].
In this work, we do not develop a privacy-preserving method but show that private information leakage is a concern for vision-language models, even when trained on anonymized data. Under this perspective, the closest work to ours is [14] which introduces a new privacy attack called Identity Inference Attack (IDIA), to assess specific individuals’ data was used in training in CLIP-like models. Differently from [14] we focus on identity recognition in generative VLMs, dealing with different critical aspects, such as prompts and context influence in recognition. Additionally, we examine the effectiveness of widely used anonymization techniques, such as face blurring, in preventing privacy leakages.
Memorization in Neural Networks
While frequently associated with over-fitting, memorization can be can happen in not over-fitted network. Specifically, studies [28, 11] hypotesized that memorization occur when the a neural network receive as learning request patterns which are harder to generalized, e.g., names. Additionaly, Duan et. al [11] show that the likelihood of a sequence being memorized increases logarithmically with its frequency in the training data and the complexity of the sequence itself. In our work, we take in account memorization pattern as a possible cause for leakage phenomena that occur even after multi-modal fine-tuning. The surprising identity recognition performance of tuned VLMs suggest strong persistency of identity association patterns, prompting future research to further investigate the correlation between memorization and data leakages.
@@ 5 Discussion
5.0.1 Pretraining and Fine-Tuning.
VLMs consist of two main components—the Language Model (LLM) and the Vision encoder. These components are pretrained on data that inherently contains personal information and identities. Although subsequent fine-tuning occurs with anonymized data, evidence suggests that the LLM and vision encoder can still memorize this sensitive information.
5.0.2 Understanding Image Embeddings.
An intriguing question arises: how does the LLM understand the image embedding of a specific person or celebrity and map it to their name? This remains an open question for future research.
5.0.3 Mitigating Privacy Risks.
How can we address the privacy risks of modern VLMs and mitigate this behaviour, one can think of different directions. a) Anonymizing Vision Encoder Output: one possibility is to anonymize the vision encoder output by adding noise or mapping the ‘private’ embedding to to generic categories (hypernyms) in order to prevent the LLM from identifying individuals. b) Unlearning Private Information: further training on both the vision encoder and the language model could help unlearn private information. However, the effectiveness of unlearning methods is not guaranteed due to the impressive memorization capabilities of large models. Never the less, additional training may impact the model’s performance on existing generic knowledge due to catastrophic forgetting, a well-studied problem in continual learning literature. c) Ad Hoc Approaches: while post-processing the LLM output seems like a straightforward solution, proxy leakage may persist. Moreover, prompt engineering can eventually break such strategies and cause leakage of private information.

### v2

@@ 1 Introduction
Our results reveal that VLMs leak identities even when modalities are aligned using anonymized datasets. This is persistent regardless of the subject’s context, showing interesting generalization capabilities. Moreover, widely adopted anonymization techniques, i.e. blurring, do not prevent privacy leakages. These results are surprising as they suggest that fine-tuning aligns unimodal (and sensible) knowledge even if not explicitly present in the training set. As a consequence, they demonstrate the need to go beyond simple data processing as a privacy protection strategies for VLMs, developing stronger strategies with more guarantees to the users. As VLMs are widely adopted, the key message of our findings is to increase our ethical awareness as a community for responsible development (and sharing) of such powerful models.
@@
2 Related Works
Vision Language Models
combine visual and textual data to solve tasks where both modalities are involved, such as Visual Question Answering [3], Image Captioning [15], Image-Text Retrieval [6] and Visual Reasoning [40]. Depending on the loss function and specific task they aim to address, VLMs are typically categorized into two main families: contrastive VLMs and generative VLMs. Contrastive VLMs are trained to assess the similarity between an image and a text. This paradigm includes widely adopted architectures like CLIP [34] and ALIGN [16] (batch-wise contrastive) and SigLIP [46] (sigmoid based).
On the other hand, generative VLMs are more powerful, as trained to generate textual descriptions or responses based on visual inputs. This family includes a diverse range of models such as BLIP [20], LLaVA [25], MiniGPT-4 [48], PaliGemma [5], and X-VLM [45]. Each of these models approaches the task with different nuances. For instance BLIP [20] leverages frozen modalities, training a fusion model. Instead, LLaVA [25] stresses the relevance of more structured data, i.e., instruction data, and removes the fusion module of BLIP [20], preferring a fine-tuning of the entire architecture. In this work, we do not propose a training paradigm or architecture but show that these architectures, may leak private information, a problem future research should account for.
Privacy-Preserving AI
seeks to protect individual data while leveraging its value for AI applications. In this direction, differential privacy [12], allows data holders to share statistical analysis while limiting what can be inferred about specific individuals by injecting calibrated noise into statistical computations. For instance, in a healthcare database, differential privacy can be applied by adding random noise to each person’s age value so that the average age of the patients can be calculated without revealing any specific individual’s age. Papernot et al [33] employs differential privacy to train models without exposing individual data points. Differently, federated learning [29] trains models on decentralized devices, keeping data local to enhance privacy. Alternatively, data anonymization protects user privacy while retaining dataset utility [36, 27, 21]. In recognition applications a simple data anonymization technique may involve blurring the face of individuals and other private information, to avoid re-identification [22].
In this work, we do not develop a privacy-preserving method but show that private information leakage is a concern for vision-language models, even when trained on anonymized data. Under this perspective, the closest work to ours is [14] which introduces a new privacy attack called Identity Inference Attack (IDIA), to assess specific individuals’ data used in training in CLIP-like models. Differently from [14] we focus on identity recognition in generative VLMs, dealing with different critical aspects, such as prompts and context influence in recognition. Additionally, we examine the effectiveness of widely used anonymization techniques, such as face blurring, in preventing privacy leakages.
Memorization in Neural Networks
While frequently associated with over-fitting, memorization may arise in not over-fitted networks too. Specifically, recent studies [28, 11] hypothesized that memorization occurs when the neural network receives as learning requests patterns that are harder to generalize, e.g., names. Additionally, Duan et. al [11] show that the likelihood of a sequence being memorized increases logarithmically with its frequency in the training data and the complexity of the sequence itself. In our work, we take into account memorization patterns as a possible cause for leakage phenomena that occur even after multi-modal fine-tuning. The surprising identity recognition performance of tuned VLMs suggests strong persistency of identity association patterns, prompting future research to further investigate the correlation between memorization and data leakages.
@@ 3.2 Architectures
As the name suggests, the objective is to embed both image and text in the same latent space, in a way that they can be compared by computing pairwise similarity. This is a core part of retrieval tasks. The matching is performed though a simple mechanism, which involves only $\boldsymbol{z}_{v}$ and $\boldsymbol{z}_{t}$. Once mapped in the same latent space, the similarity between the text and image can be estimated using any metric distance $s(\boldsymbol{z}_{v},\boldsymbol{z}_{v})$.
@@ 4.1 Experimental setting
Figure 3: Example of the picture manipulation we evaluated. (a) Original picture. (b) Background replaced with a random landscape background. (c) Background replaced with white to force the model focus towards the subject. (d) Face blur to question its effectiveness in preventing leakages.
The first step is to collect a dataset of people whose images are contained in pre-training datasets, e.g., LAION [37]. The choice goes to publicly available pictures of celebrities, following a similar approach of [34] and [14]. We collected pictures of celebrities from publicly available sources, as no datasets with identity-face association are currently openly available. We selected 500 (Appendix  0.D) celebrities and collected 50 pictures each, summing up to 25k images for our experiments.
To remove potential sources of errors, we excluded pictures containing written names, which could guide VLMs to correctly recognize the person only from the OCR capability. To do so we relied on EasyOCR [1] to detect text in the pictures, excluding only those containing a sub-string of the name with a minimum length of 5 characters. In total, only 1171 images have been removed.
@@
Generative VLMs are sensitive to the specific prompt used to condition their outputs. To avoid our findings being limited by specific prompt choices, we studied models’ behavior when conditioned on different prompt categories, and their rephrasing. We identified five prompts representing increasing levels of details required from the model:
@@
Interestingly, while recent research has analyzed identity recognition using contrastive VLMs [14], the privacy risks and the memorization of individuals’ identities remain unexplored in the context of generative VLMs. We opt for 5 VLMs pre-trained on potentially private data [37] and fine-tuned with few to no private data:
@@ 4.2 Do VLMs leak names?
Model $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$leaks ±$\pm~{}\sigma$ BLIP-2flan-t5-xl [20] $0.948\pm$0.117 $1.011\pm$0.880 $6.943\pm$0.017 $4.494\pm$6.196 $7.208\pm$0.004 BLIP-2opt [20] $1.428\pm$0.487 $2.847\pm$3.410 $5.235\pm$3.747 $3.471\pm$2.999 $4.603\pm$4.994 LLaVA-1.57B [24] $0.124\pm$0.001 $0.062\pm$0.003 $4.139\pm$8.991 $3.593\pm$9.692 $9.557\pm$0.311 LLaVA-1.6mixtral-7B [24] $0.352\pm$0.053 $0.182\pm$0.001 $7.120\pm$1.742 $1.496\pm$2.233 $6.827\pm$1.023 PaliGemma3b-mix-224 [5] $1.116\pm$1.978 $2.787\pm$2.573 $7.808\pm$0.729 $6.184\pm$2.079 $7.659\pm$0.700
@@
Looking at the results, we see that name leakages (or identity recognition) rate is not negligible, with a peak rate of 9% ($\sim$2250) of pictures whose subject has been identified by LLaVA-1.57B [24]. Second, all models leak more when directly asked to generate a person name, but can still leak correct identities even when prompted to generically describe the picture. For instance, the average leakage rate under $P_{2}$ (“Who is the person in the picture”) is $\sim$6%, with a peak of 7.81% with PaliGemma3b-mix-224 [5], while with $P_{3}$ (“Describe the celebrity in the picture.”), the average leakage rate is 3.4%. Looking at the most generic prompt $P_{0}$ (“Describe the picture.”), leakages decrease, but their rate is still not negligible with BLIP-2flan-t5-xl [20] (1.43% or 357 leaked images) and with PaliGemma3b-mix-224 [5] (1.12% or 280 leakaged images). This not only implies that identity recognition can be a threat, but that it could happen in an unpredictable way, i.e., even when the model is not explicitly asked to perform it.
We underline that this happens even when models are fine-tuned on anonymized data. In fact we do not see any significant reduction in leakage for VLMs finetuned with anonymized data compared to those finetuned with partially anonymized ones (BLIP-2). This indicates that data anonymization does not address leakage if parts of the model (e.g., visual/textual backbones) are already exposed to private information. Additionally, some models (i.e., LLaVA-1.57B [24] and BLIP-2opt [20]) are more susceptible to prompt rephrasing, as the high standard deviations show. In the next experiments we use only one prompt per category, i.e., the 5 prompts reported in Sec. 4.1.2, as their leakage rate is the closest to the average leakage rate per category (see Tab. 7 for the expanded results).
@@ 4.3 How does background influence leakages?
For instance, it is easier to identify Cristiano Ronaldo in a football field, rather than Cristiano Ronaldo alone, given the positive correlation between the two elements. To disentangle this correlation and focus primarily on the subject, we isolate the celebrities in each picture and change the background. To do so we used Grounding Dino [26] to first detect the class “person” in the picture, identifying their bounding boxes. After this, we forwarded each box to SAM [19] to obtain each subject’s segmentation masks. Finally, we extracted the person’s masks and blend it with different backgrounds as shown in Figure 3 (b) and (c). We consider two cases of background choice.
@@
Results in Table 3 confirms that the reducing context distractions (as in landscape backgrounds) leads to higher leakages compared to cases where the background is a scene. For instance, LLaVA-1.6mixtral-7B [24] goes from 5.18% on $P_{4}$ with landscape background to 6.41% with white background. This highlights that landscape background can introduce novel patterns that weaken model’s focus on the subject, degrading their performance in recognizing identities. However, the overall performance remain in the same range, demonstrating that in both settings the model is able to focus on the subject and recognize their identity. Again, we underline how robust identity recognition is, as not only all the models are able to recognize people from images that could have been precedently memorized, but instead they are able to recall the identity association even in different contexts.
@@ 4.4 Can blurring prevent privacy leakages?
This outcome suggests that blurring, without compromising the content of an image, may not be an adequate technique for anonymization. A malicious attacker could use, for instance, Google Maps to recognize people whose pictures have been inadvertently included in a pretraining dataset.
4.4.2 Blurring on the replaced background.
@@ 4.5 Statistical analysis on leaked celebrities and discussion
To further complement the experiments, we conduct a statistical analysis of the leaked celebrities, trying to extract insights that may suggest the origin of this behavior.
@@
We computed the leakage rate per celebrity (Figure 4), showing that more famous celebrities seem to be identified with more ease. Then we plot the distribution of the leakages per celebrity 5(a), which highlights that while a discrete portion of celebrities have been rarely identified, another portion of them have been easily identified throughout all the settings.
@@
As major continual learning works [23, 2, 43] emphasize the impact of catastrophic forgetting after additional sessions of training, face-identity recognition after fine-tuning on anonymized data is surprising. This may be partially explained by the phenomenon known as memorization [11, 28]. When abnormal or complex patterns, such as the names of people, are presented to a model, it tends to memorize these specific instances rather than broadly learning general concepts. This phenomenon suggests that face-identity associations might be stored within a few neurons of the model as distinct activation patterns. These stored patterns can potentially be retrieved with some degree of context generalization, meaning that even slight contextual clues could trigger the recall of specific identities.
This behavior raises significant privacy concerns, as it implies that sensitive information could be inadvertently exposed. Therefore, we need future research aimed at preventing such privacy leakages from occurring in generative Visual-Language Models (VLMs). Researchers will need to develop methods to ensure that models learn generalizable concepts without memorizing sensitive or personal information, thereby preserving users’ privacy.
@@ 5 Discussion
Pretraining and Fine-Tuning. VLMs consist of two main components—the Language Model (LLM) and the Vision encoder. These components are pretrained on data that inherently contains personal information and identities. Although subsequent fine-tuning occurs with anonymized data, evidence suggests that the LLM and vision encoder can still memorize this sensitive information.
Understanding Image Embeddings. An intriguing question arises: how does the LLM understand the image embedding of a specific person or celebrity and map it to their name? This remains an open question for future research.
Mitigating Privacy Risks. To address the privacy risks of modern VLMs, several approaches can be considered: a) Anonymizing Vision Encoder Output: This involves adding noise or mapping private embeddings to generic categories, preventing the LLM from identifying individuals. b) Unlearning Private Information: Additional training on the vision encoder and language model might help unlearn private information, though its effectiveness is uncertain due to the models’ strong memorization capabilities. Additionally, this could also harm performance due to catastrophic forgetting. c) Ad Hoc Approaches: Post-processing LLM output or using prompt engineering may help, but these methods may still allow proxy leakage of private information.